OpenStudy (anonymous):

Rewrite the expression using positive exponents -3^0/4x^-3

OpenStudy (anonymous):

X ^ 0 = 1 X ^ -y = 1/X^y Got it?

OpenStudy (anonymous):

i really didnt quit get it i get the part that x to the zero is 1

OpenStudy (anonymous):

now for the second part: 4x^(-3) = 1/4x^(3)

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

can you tell me what the final answer would be

OpenStudy (anonymous):

final answer: 4x^3

OpenStudy (anonymous):

so my answer would not be 1/4x^3